Daily Report

CPUC/ANF Environmental Monitoring Activities

Pacific Pipeline Construction Project

Aspen Environmental Group Thursday 10, December 09, 1998

The following is a brief summary of the environmental monitoring activities that were completed on December 10, 1998. This summary is based on reports directly received from the field and has not been edited. If you have any questions please contact (818) 597-3407. Please note the item number and the correction that should be made. This report is intended to serve as a communication document and as an instrument for quality control.

Construction Highlights

12/09/98 (Wednesday)

Segment 1 report for 12/9/98: At Emidio Pump Station, commissioning of the pump equipment and work on the office interior continued. Clean-up was underway on Hwy. 166. A segment of trench (EMP 8.95 to 8.92) is being excavated so pipe can be placed and tied-in to the pipe through Stream Crossing #4 (EMP 9.03). Negative buoyance pipe was being delivered to the California Aqueduct bore. Bore operations continued at Stream Crossing #9 (EMP 15.35). On Grapevine Rd., welding and x-ray occurred from 15.56 to 15.52. Shading took place from 16.20 to 16.02. At Grapevine Pump station, electrical work continued and forming of the platform around the substation was under way. On Tejon Bench, trenching for a segment of pipe that was skipped was underway at EMP 16.87. Backfill was underway to EMP 16.64. Excavation occurred to EMP 18.08 (working north) and backfill has been done on that section to EMP 18.12. On Lebec Rd. backfill occurred from EMP 21.46 to 21.50 (working south). A segment of pipe was being tied-in at 21.54 and trench has been excavated to 21.56. Pipe is being pushed through the bore at the Cuddy Creek Tributary, EMP 25.47.

Segment 2B (Castaic to Sylmar): Report for 12-9-98. Slurry backfill and fiber optic duct were laid south of Weldon Cyn near EMP 72.2. In addition, paving and clean up progressed between Sierra Hwy and Weldon Cyn.

Segment 3A & 3C (Sylmar to Los Angeles): Report for 9 December 1998: 3A: NIGHT work continued at San Fernando and Hollywood Way Bridges. The crews completed installation of insulation at San Fernando and grouted the south end, around the sleeve. Work will continue tonight. Welders, x-ray, and coating crews spent most of the day working on the bore tie in under Tuxford Street. X-ray, coating, and backfill crews had a short day working on the Pacoima Wash Bridge crossing. The crews were shut down by MTA due to the need for the flagmen to attend a company function. Minor backfill and compaction of sink holes and assist pits were done along the right of way in Glendale and Burbank. 3C: Crews placed a fiber optic pull box near the aqueduct crossing. Crews working on the high speed channel had to stop early due to the high winds.

Segment 4B (Lynwood to EWl Segundo): Report for 12-9-98. At the Lynwood station, paving crews repaired the ditch along little Alameda St while compaction and wall building continued at the station. In addition, crews progressed with final electrical tie-ins on the valves. Crews working at the El Segundo station completed welds and began preparations for hydro-testing. Further, clean up crews continued work along the rail road right of way. Segment 5: At the Wilmington station (Tosco) crews removed metering turbines that rusted after exposure to water. In addition, crews are completing electrical work on the station house and valve bodies and hydro-test crews welded a test head on the pipe running to the station from Alameda St. Fiber optic cable was pulled from Del Amo Blvd to the 91 Freeway while a sideboom was removed from the right of way.

Inspection Summary

In the past 24 hours, 4 inspection event(s) were reported for 4 spread(s). Of these inspections, 0 were non-compliance event(s). Also 0 non-compliance resolution report(s) was resolved.

Non-Compliance Reports (NCR's) Issued

No new non-compliance reports were issued during the last reporting period.

Non-Compliance Resolution Reports (NCRR's) Issued

Spread 1: NCRR Issued

NCRR# 1110332 Cleared NCR#110008 Date NCR Issued: 12/10/98 Monitor's Name: Etchell
Inspection Code: 3 Location:  Emidio to ANF
Begin Station: 0 End Station: 33 Landowner: Other Activity: General

Summary of Follow-up Items for Today

No follow-up items were reported for today.

Back to Daily Reports

Back to MMCRP Home Page